# Single-Channel EEG Tokenization Through Time-Frequency Modeling

### v1

@@ A.4. Ablation on Token Vocabulary Size
Figure 7. Token vocabulary size ablation with performance metrics
Figure 8. Token vocabulary size ablation with token utilization and uniqueness
A.5. Ablation on Masking
Table 7. Ablation on masking used for the pretraining of TFM-Tokenizer on TUEV Dataset
@@ A.5. Ablation on Masking
Appendix B TFM-Token Implementation and Hyperparameter Tuning
Figure 9. TFM-Token Overview
Figure 9 presents an overview of TFM-Token during inference. This section provides additional details on the implementation and training of the framework.
B.1. Training Pipeline:
For all experiments, we follow a single-dataset setting, where all processes in each experiment are conducted within the same dataset. The training process of our framework is as follows: (1) TFM-Tokenizer unsupervised pretraining, (2) TFM-Encoder pretraining using masked token prediction, and finally (3) fine-tuning on the same dataset for downstream tasks.
B.2. Hyperparameter Tuning of TFM-Token
We employed a systematic approach to optimize the hyperparameters of both the TFM-Tokenizer and TFM-Encoder models using Ray Tune [^1] https://docs.ray.io/en/latest/tune/ with the Optuna [^2] https://optuna.org/ search algorithm. Our optimization process followed a three-phase strategy.
In the first phase, we optimized the TFM-Tokenizer architecture by tuning the depth and number of attention heads in the frequency transformer, temporal transformer, and transformer decoder modules to minimize the masked reconstruction loss $\mathcal{L}_{recon}$. This was followed by tuning the training optimizer’s parameters, including learning rate and weight decay. The second phase focused on the TFM-Encoder optimization for the classification task, where we first tuned its architectural parameters (depth and number of heads), followed by training the optimizer’s parameters while keeping the tokenizer frozen. The third phase focused on tuning optimizer parameters for the masked token prediction pretraining of the TFM-Encoder.
To ensure a fair comparison with LaBraM’s neural tokenizer, we maintained a vocabulary size of $8,192$ and an embedding dimension of $64$. For our ablation studies involving raw signal-only and STFT-only variants, we doubled the embedding dimensions of the temporal encoder and frequency patch encoder to match the codebook dimension while maintaining all other parameters same. Detailed hyperparameter configurations for both TFM-Tokenizer and TFM-Encoder are provided in Appendices B.3 and B.4, respectively.
B.3. TFM-Tokenizer Hyperparameters
Table 8. Hyperparameters for TFM-Tokenizer unsupervised pretraining on single-channel setting
@@ B.3. TFM-Tokenizer Hyperparameters
Table 9. Hyperparameters for TFM-Tokenizer
@@
B.4. TFM-Encoder Hyperparameters
Table 10. Hyperparameters for TFM-Encoder, its masked token prediction pretraining and downstream finetuning
@@ B.4. TFM-Encoder Hyperparameters
Appendix C Additional Experiment Details
C.1. Dataset Statistics and Splits
Table 11. Dataset Summary
Dataset # of Recordings # of Samples Duration (s) Task
TUEV $11,914$ $112,491$ $5$ EEG Event Classification
IIIC Seizure $2,689$ $135,096$ $10$ Seizure Type Classification
CHB-MIT $686$ $326,993$ $10$ Seizure Detection
TUAB $2,339$ $409,455$ $10$ Abnoral EEG Detection
This section provides detailed information on the datasets used in our experiments and their respective splits. Table 11 summarizes key statistics, including the number of recordings, the total number of samples after preprocessing, their duration, and the corresponding downstream tasks. For TUEV and TUAB, we utilized the official training and test splits provided by the dataset and further divided the training splits into $80\%$ training and $20\%$ validation sets. We performed a subject-wise split into $60\%$ training, $20\%$ validation, and $20\%$ test on the IIIC Seizure dataset. In the CHB-MIT dataset, we used 1-19 subjects for training, 20-21 for validation, and 22-23 for testing.
C.2. STFT parameters
To extract frequency-domain representations of the EEG, we utilized the STFT function from PyTorch. The recommendations of (Yang et al., 2024) guided our parameter selection and empirical analysis of different configurations to optimize the time-frequency resolution tradeoff. The final parameters are as follows:
Table 12. STFT parameters
Parameter Value Description
Hop length $H$ $100$ Step size for sliding window ($50\%$ overlap)
Centering False The STFT is computed without implicit zero-padding

### v2

@@ C.8 Ablation on Token Vocabulary Size
Figure 9: Token vocabulary size ablation with performance metrics
Figure 10: Token vocabulary size ablation with token utilization and uniqueness
C.9 Ablation on Masking
Table 12: Ablation on masking used for the pretraining of TFM-Tokenizer on TUEV Dataset
@@ C.9 Ablation on Masking
Appendix D TFM-Tokenizer Implementation and Hyperparameter Tuning
Figure 11: TFM-Tokenizer framework Overview
Figure 11 presents an overview of the framework during inference. This section provides additional details on the implementation and training of the framework.
D.1 Hyperparameter Tuning of TFM-Tokenizer and Downstream Transformer
We employed a systematic approach to optimize the hyperparameters of both the TFM-Tokenizer and downstream transformer models using Ray Tune [^1] https://docs.ray.io/en/latest/tune/ with the Optuna [^2] https://optuna.org/ search algorithm. Our optimization process followed a three-phase strategy.
In the first phase, we optimized the TFM-Tokenizer architecture by tuning the depth and number of attention heads in the frequency transformer, temporal transformer, and transformer decoder modules to minimize the masked reconstruction loss $\mathcal{L}_{recon}$. This was followed by tuning the training optimizer’s parameters, including learning rate and weight decay. The second phase focused on the downstream transformer optimization for the classification task, where we first tuned its architectural parameters (depth and number of heads), followed by training the optimizer’s parameters while keeping the tokenizer frozen. The third phase focused on tuning optimizer parameters for the masked token prediction pretraining of the downstream transformer.
To ensure a fair comparison with LaBraM’s neural tokenizer, we maintained a vocabulary size of $8,192$ and an embedding dimension of $64$. For our ablation studies involving raw signal-only and STFT-only variants, we doubled the embedding dimensions of the temporal encoder and frequency patch encoder to match the codebook dimension while maintaining all other parameters same. Detailed hyperparameter configurations for both TFM-Tokenizer and downstream transformer are provided in Appendices D.2 and D.3, respectively.
D.2 TFM-Tokenizer Hyperparameters
Table 13: Hyperparameters for TFM-Tokenizer unsupervised pretraining on single-channel setting
@@ D.2 TFM-Tokenizer Hyperparameters
Table 14: Hyperparameters for TFM-Tokenizer
@@
D.3 Downstream Transformer Encoder Hyperparameters
Table 15: Hyperparameters for downstream transformer, its masked token prediction pretraining and downstream finetuning
@@ D.3 Downstream Transformer Encoder Hyperparameters
Appendix E More Related Works
Frequency Representation Collapse. Frequency domain analysis is crucial in EEG and general time series analysis  (Elvander & Jakobsson, 2020; Wu et al., 2021; 2023; Woo et al., 2022). In real-world signals, time-domain observations inherently mix multiple frequency components, and high-energy, low-frequency signals often dominate the spectrum  (Huang Norden E Shen Zheng & H, 1998; Lai et al., 2018). As a result, these entangled frequency features makes it difficult for models to distinguish between them  (Zhou et al., 2022; Piao et al., 2024). Recent studies have shown that these entangled signals can lead to a collapse in the learned frequency representations  (Zhi-Qin John Xu et al., 2020; Piao et al., 2024). Models tend to overemphasize the dominant low-frequency features while neglecting the high-frequency details. This issue can lead to a lack of capturing various EEG waveforms and degenerating data representation  (Park & Kim, 2022). Motivated by these works, our paper focuses on developing methods to learn diverse, informative frequency features. In Section 7, we provide an analysis of our proposed frequency-domain tokenizer and its impact on model performance.
Appendix F LLM Usage Statement
We used large language models (LLMs) solely for writing support, including grammar correction, sentence refinement, and clarity improvements. All conceptual contributions, algorithm design, code development, experiments, and analyses were conducted entirely by the authors.